EXPERIMENTAL ROTATIONAL SPECTRA FOR MnRe(CO)_{10} AND o-C_{6}H_{4} REQUIRED ACCURATE THEORETICAL CALCULATIONS FOR SUCCESSFUL ANALYSIS^{a}

\underline{STEPHEN G. KUKOLICH}, Department of Chemistry, University of Arizona, Tucson, AZ 85721; MICHAEL H. PALMER, School of Chemistry, University of Edinburgh, Edinburgh EH9 3JJ, United Kingdom; PETER GRONER, Department of Chemistry, University of Missouri-Kansas City, Kansas City, MO 64110; and CHAKREE TANJAROON, Department of Chemistry, University of Alberta, Edmonton, Alberta T6G 2G2, Canada.

The keys to successfully analysing the high-resolution rotational spectra for manganese-rhenium decacarbonyl and for o-benzyne were the results of extensive and accurate theoretical calculations. Obtaining the microwave spectrum of the dinuclear complex MnRe(CO)_{10} was made difficult by the low rotational constant (200 MHz) resulting in fairly high J values for transitions in the normal Flygare-Balle spectrometer region. The assignments were further complicated by the presence of ^{185}Re, ^{187}Re isotopes with eQq = 391 and 370 MHz and ^{55}Mn with eQq = -16.5 MHz. Large Gaussian orbital basis calculations by Palmer and Guest provided sufficiently accurate predicted molecular constants to assign and fit the spectrum. Although the spectrum for the normal isotopomer of o-benzyne could be obtained with reasonable S/N, assigning the singlesubstitution ^{13}C isotopomer spectra, even with an enriched sample, proved to be difficult. The accurate and precise rotational constants for the 7 measured isotopomers for o-C_{6}H_{4} could not be accurately fit using a planar structure due to the vibrational averaging effects. Using ab initio vibrational averaging corrections calculated by Groner, an accurate, near-equilibrium structure was obtained for o-benzyne.

aSupported by THE NATIONAL SCIENCE FOUNDATION